### v1

@@ Abstract
Meeting net-zero targets remains formidable as terrestrial renewables grapple with intermittency and regional variability. Here, we integrate space-based solar power (SBSP)—a potential near-constant, orbital solar technology—into a high-resolution, Europe-wide capacity-expansion and dispatch model to quantify its contribution under net-zero constraints. We examine two advanced SBSP designs: (1) a near-baseload, low Technology Readiness Level (TRL) concept (heliostat-based Representative Design RD1) and (2) a partially intermittent, higher-TRL concept (planar-based RD2), both drawing on NASA’s 2050 cost and performance projections. Our results show that RD1 can reduce total system costs by 7–15%, displace up to 80% of intermittent wind and solar, and cut battery usage by over 70%, if it meets its forecast cost reductions—though long-duration storage (e.g., hydrogen) remains essential for seasonal balancing. By contrast, RD2 is economically unattractive at its projected 2050 costs. Through extensive sensitivity analyses, we identify cost thresholds at which SBSP shifts from cost-prohibitive to complementary and ultimately to a dominant baseload technology. Specifically, RD1 becomes complementary at roughly 14× and dominant at 9× the 2050 solar PV capital cost, benefiting from its continuous power generation. Meanwhile, RD2 must achieve even lower cost levels (9× to be complementary and 6× to dominate) and would rely on short-duration storage to mitigate its partial intermittency. These findings provide quantified techno-economic benchmarks and reveal alternative net-zero pathways, offering critical guidance for policymakers and industry stakeholders seeking large-scale, centrally coordinated renewable solutions with non- or low-intermittency.
@@ 1 Background
Space-based solar power (SBSP) could offer a centralized, weather-independent energy resource to help address these challenges [6]. By operating above the atmosphere and outside the day-night cycle, SBSP promises continuous gigawatt-scale power. Recent technological milestones suggest it may evolve from a niche concept to a techncially viable solution by the 2030s: multi-junction and lightweight PV cells achieving near 47% efficiency [7, 8], modular in-orbit assembly [9], and successful wireless power demonstrations [10, 11] have all reached mid-range Technology Readiness Levels (TRL 4–5). Meanwhile, reusable launch vehicles have significantly reduced launch costs, with some estimates projecting a Levelized Cost of Energy (LCOE) for SBSP of \$30–80 MWh-1 by 2050 [12, 13, 14, 15]. Although considerable uncertainties remain—from in-orbit manufacturing to policy frameworks—major space agencies (NASA, ESA, JAXA) are actively shaping regulatory pathways, motivating the need to understand SBSP’s potential contribution to net-zero goals [14, 16, 17].
@@
  • Integrate SBSP into a continental-scale model: We embed SBSP into a European-scale capacity-expansion and dispatch framework, allowing us to evaluate its interactions with terrestrial renewables, storage, and transmission infrastructure under net-zero constraints.
  • Assess grid balancing benefits: We quantify how more continuous SBSP can reduce storage requirements and transmission investments by offsetting the temporal and spatial variability of terrestrial wind and solar.
@@ 2.1 SBSP Concept
Figure 1 illustrates the main stages of an SBSP system, from satellite deployment in orbit to grid integration on Earth. Most designs target geostationary Earth orbit (GEO) because the orbital period matches Earth’s rotation, minimizing eclipse times to under 1% of operation [18]. This near-continuous solar exposure allows stable, high-capacity power generation.
After launch, specialized robotic systems assemble modular components in orbit, including lightweight photovoltaic (PV) arrays or concentrating mirrors for solar collection. Two principal approaches are used: (i) reflectors that concentrate solar radiation onto a central receiver, and (ii) direct solar absorption via PV panels [19, 20]. These methods convert solar irradiation into electrical power, forming the core energy-harvesting mechanism.
The harvested electricity is then converted into microwave or laser beams for transmission. Microwave frequencies (1–10 GHz), especially 2.45 GHz or 5.8 GHz [21], are commonly selected to balance transmission efficiency, atmospheric attenuation, and safety constraints [22, 23]. Although lasers offer higher power densities, they are generally limited to low Earth orbits due to beam divergence and safety considerations. In this study, microwave beams are considered.
@@
On the ground, a rectifying antenna (rectenna) spanning several square kilometers captures the transmitted energy and converts it to direct current (DC). This DC power is subsequently inverted to alternating current (AC) and fed into the existing grid infrastructure. Because rectennas employ lightweight mesh designs, they allow partial land co-use. Multiple ground stations can also receive power from a single SBSP satellite, facilitating broader distribution [18].
Finally, standard transformers and transmission lines deliver the electricity to end users. A dedicated control center manages satellite pointing, power beaming, and system diagnostics, ensuring real-time coordination between orbit and ground. This centralized architecture, leveraging continuous solar availability in GEO, offers a potentially high-capacity, dispatchable renewable option for meeting baseload and balancing demands.
Figure 1: Operational Steps of a SBSP System. (1) Launch and installation of the SBSP system in space; (2) Collection of solar energy in space; (3) Conversion of solar energy into electricity in space; (4) Transmission of microwave to ground stations on Earth; (5) Ground stations receive and convert the transmitted energy into electricity; (6) Delivery of electricity to the grid and users.
Figure 2: The main structures of The Innovative Heliostat Swarm and The Mature Planar Array systems. (a) The Innovative Heliostat Swarm design employs a beehive concept, where hexagonal small components operate independently to form a large satellite. (b) The Mature Planar Array model consists of a large panel with a capability of power generation/transmission and a bus system which are connected by multi-wires. It uses a “sandwich” architecture, where solar energy is collected on one side and coherent RF is transmitted from the other. The power generation/transmission panel consists of numerous identical power modules, each converting solar energy to microwave power. Modules are wirelessly controlled by the bus system, eliminating the need for wired signal or power connections. To maximize efficiency, the microwave antennas are kept aligned toward Earth by gravity gradient forces in a GEO.
@@
Although originally conceived by Peter Glaser in 1968 [24], SBSP remained impractical in early evaluations due to prohibitively high costs for launch, in-orbit assembly, and maintenance [25, 26]. Over the past three decades, however, significant advances in solar cell efficiency, wireless power transmission, and reusable launch vehicles have reignited interest in SBSP. A variety of novel designs have since emerged, including the Large Array of Solar Panels [24], Sun Tower Design [26], Modular Symmetrical Concentrator [27], Solar Power Satellite via Arbitrarily Large Phased Array [28], Space-Based Reflectors [29], Inflatable Thin-Film Structures [30], Hypermodular SBSP Design [31], and Tethered Satellite System [32]. Each concept seeks to optimize power density, structural mass, and reliability under the demanding conditions of orbit.
@@
  • Representative Design One (RD1), the Innovative Heliostat Swarm: Based on the SPS-ALPHA Mark III concept [34, 35, 36], this lower-TRL design uses mirror-like reflectors (heliostats) to direct sunlight to a central concentrator, enabling nearly 99.7% annual power availability [34]. Key innovations include retro-directive RF transmission arrays, high-efficiency PV cells, lightweight modular structures, and autonomous in-orbit deployment. Figure 2(a) shows how reflectors continuously adjust orientation, maximizing solar capture throughout the orbital cycle.
  • Representative Design Two (RD2), the Mature Planar Array: Adapted from JAXA’s Tethered Solar Power Satellite [20] and Caltech’s SSPP design [37], RD2 includes planar panels whose lower surface faces Earth under gravity-gradient forces. While solar incidence on the upper and lower surfaces varies with orbital geometry, the system achieves roughly 60% annual power availability [20]. Figure 2(b) illustrates its higher TRL, given demonstrated hardware and well-documented performance characteristics.
These two concepts differ in TRL and technical approach yet both incorporate high-efficiency solar capture, wireless power transmission, and modular architectures designed to lower launch mass and total costs. By examining RD1 (near-baseload, lower TRL) and RD2 (planar, higher TRL) side-by-side, we provide a balanced assessment of SBSP performance, feasibility, and cost. This approach captures essential trade-offs in orbital assembly complexity, antenna design, and seasonal power availability, ensuring our analysis thoroughly reflects the technical and economic considerations that will guide future SBSP deployment.
@@ 2.2 SBSP Modeling
To assess the system-level impacts of SBSP, we modeled the generation profiles and lifetime costs of NASA’s two representative designs (RD1 and RD2) for both current (2020) and future (2050) scenarios. Drawing on published technical data [33], we derived hourly generation outputs by accounting for orbital geometry, eclipse periods, solar incidence angles, and design-specific operational constraints (e.g., RD1’s heliostat-based approach vs. RD2’s planar configuration). We assume the geostationary orbit altitude ($\sim 35{,}786\text{ km}$) is negligible relative to the Earth–Sun distance ($\sim 1.49\times 10^{8}\text{ km}$) [38, 33]. Complete derivations and assumptions, including the inverse-square law for irradiance and eclipse scheduling, are detailed in the Supplementary Information (SI, Section 1).
Generation Profiles. Both RD1 and RD2 capture solar energy in GEO nearly year-round, but differ in achievable annual availability: RD1 uses heliostats to concentrate sunlight, enabling up to 99.7% power availability, while RD2’s planar panels achieve approximately 60% due to geometric constraints [34, 20]. Compared to terrestrail solar panels whose power avaialbility is usually 15-30% [39], both RD1 and RD2 show improved power availability. We simulate and compare the hour-by-hour power output per square meter of RD1, RD2, and terrestrial solar panels, based on representative orbital parameters, solar irradiation, panel characteristics, and efficiencies (SI, Section 1). Using 2020 as a case study, the results shown in Figure 3 demonstrate the potential advantages of SBSP power generation. The power availability of RD1 and RD2 exceeds that of terrestrial solar by 70% and 42%, respectively. Furthermore, their power output is more continuous, significantly improving efficiency, especially for RD1. During the continuous phase, the power output per square meter of RD1 and RD2 is higher at different times compared to terrestrial solar. Over the course of 2020, the power output per square meter of RD1 and RD2 panels is 384% and 233% higher than that of terrestrial solar, respectively. Terrestrial solar has higher output in summer than in winter, while SBSP exhibits the opposite trend, providing complementary power generation in the future. Additionally, the power output of SBSP is significantly less affected by seasonal changes compared to terrestrial solar, indicating its potential for more stable power generation with smaller fluctuations.
@@
Figure 3: Comparison of the variations in power output per square meter of RD1, RD2, and terrestrial solar panels in 2020. (a) To clarify the power generation curves of RD1, RD2, and terrestrial solar, the hourly data were averaged to obtain daily mean values, enabling the observation of overall trends throughout 2020. (b) Hourly data from July 1 to July 15 represent the summer period. (c) Hourly data from December 1 to December 15 represent the winter period.
Cost Analysis. We adopt NASA’s five-phase concept of operations (ConOps) framework (Develop, Assemble, Operate, Maintain, and Dispose) to estimate capital expenditures (CapEx) and fixed operation and maintenance costs (FOM) [33]. The Develop and Assemble phases, covering technology R&D, hardware manufacturing, launch, and in-orbit assembly, are aggregated into CapEx, while Operate, Maintain, and Dispose phases contribute to FOM. Key cost drivers include reusable launch vehicles, modular assembly robotics, and lifetime maintenance. Variable Operations and Maintenance (VOM) costs typically apply to electricity fuel sources. However, since SBSP does not rely on fuel as a direct input for electricity production, this cost category, along with Fuel Costs, is excluded for all renewable electricity production technologies [33]. Under baseline assumptions, we find that near-term (2020) SBSP costs are high, but can drop by 2050 with NASA’s projected efficiencies, launch cost reductions, and technological improvements (SI, Section 2). Building on this foundation, we conducted multiple variable sensitivity analyses, incorporating predicted data for various indicators based on NASA’s projections for 2050 [33]. The estimated costs of RD1 and RD2 in 2020 (the baseline case) and 2050 are shown in Table 1. CapEx refers to the capital expenditures required to acquire, upgrade, or maintain long-term assets, such as equipment, while FOM represents the annual costs associated with the fixed operation and maintenance of these assets. VOM, on the other hand, refers to the variable operation and maintenance costs that change with energy production or system usage [40].
@@
Table 1: Key Parameters and Costs for SBSP under baseline and 2050 scenarios.
@@
RD1 Baseline 30 2,028.79 99.7 3
RD2 Baseline 30 2,021.95 60 3
RD1 2050 30 2,028.79 99.7 3
RD2 2050 30 2,021.95 60 3
CapEx (\$/kW) FOM (\$/kW-year) VOM (\$/kWh) Fuel Cost (\$/MMBtu)
44,642.34 3,050.26 N/A N/A
67,756.43 4,899.07 N/A N/A
3,168.38 129.51 N/A N/A
4,024.02 225.77 N/A N/A
Table 2: Cost data for SBSP in models (PyPSA-Eur) under baseline and 2050 scenarios [33].
SBSP Capital cost (EUR/kW) Marginal cost (EUR/kW)
RD1 Baseline 4,901.65 0
RD2 Baseline 7,687.48 0
RD1 2050 267.87 0
RD2 2050 396.59 0
@@ 3 Results
Table 3 provides an overview of the key characteristics of the scenarios analyzed in this study. These scenarios encompass variations in network configurations, technology cost assumptions (generators, storage, and transmission) and the inclusion or exclusion of SBSP with different designs (RD1 and RD2). Scenario 1 is designed for the assessment of SBSP under the 2020 baseline. Scenarios 2–10 focus on assessing SBSP under predicted 2050 conditions (based on the cost data in Table 2). Scenarios 11–26 are dedicated to sensitivity analyses, exploring the impact of varying costs of SBSP (analyzed in the sensitivity analysis).
The SBSP-2020 scenario - Scenario 1
Scenario 1 represents a current-cost baseline for assessing the cost-effectiveness of SBSP in today’s energy system. We individually integrate RD1 and RD2 into the PyPSA-Eur model for minimum-cost optimization (see “Methods”). In this 2020 context, RD1 (4900 EUR/kW) and RD2 (7690 EUR/kW) are approximately 98 and 154 times more expensive than terrestrial solar (50 EUR/kW), respectively, and also greatly exceed onshore wind (110 EUR/kW), reservoir hydro (220 EUR/kW), biomass (240 EUR/kW), offshore wind (280 EUR/kW), run-of-river hydro (310 EUR/kW), and nuclear (580 EUR/kW). Their capital costs are thus orders of magnitude higher than other technologies, rendering SBSP economically non-competitive in 2020 (Fig. 4). These results further underscore the current economic infeasibility of SBSP. Despite being integrated, the optimization results reveal that neither RD1 nor RD2 is utilized under the given conditions, with both designs showing zero installed capacity. This finding suggests that SBSP is currently unviable for deployment under existing conditions due to its high capital cost.
Table 3: Descriptions of different scenarios
Scenario Description Scenario Description
1 The entire network configuration is derived from data corresponding to the year 2020. 14 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 300 EUR/kW.
2 A predicted 2050 network where all technology costs are set to their forecasted maximum values, and SBSP is not included in the network. 15 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 250 EUR/kW.
3 A predicted 2050 network where all technology costs are set to their forecasted maximum values, and SBSP is included in the network with the RD1 design. 16 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 200 EUR/kW.
4 A predicted 2050 network where all technology costs are set to their forecasted maximum values, and SBSP is included in the network with the RD2 design. 17 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 150 EUR/kW.
5 A predicted 2050 network where all technology costs are set to their forecasted average values, and SBSP is not included in the network. 18 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 100 EUR/kW.
6 A predicted 2050 network where all technology costs are set to their forecasted average values, and SBSP is included in the network with the RD1 design. 19 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 450 EUR/kW.
7 A predicted 2050 network where all technology costs are set to their forecasted average values, and SBSP is included in the network with the RD2 design. 20 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 400 EUR/kW.
8 A predicted 2050 network where all technology costs are set to their forecasted minimum values, and SBSP is not included in the network. 21 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 350 EUR/kW.
9 A predicted 2050 network where all technology costs are set to their forecasted minimum values, and SBSP is included in the network with the RD1 design. 22 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 300 EUR/kW.
10 A predicted 2050 network where all technology costs are set to their forecasted minimum values, and SBSP is included in the network with the RD2 design. 23 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 250 EUR/kW.
11 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 450 EUR/kW. 24 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 200 EUR/kW.
12 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 400 EUR/kW. 25 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 150 EUR/kW.
13 Based on Scenario 6, with all other parameters unchanged, the capital cost of RD1 is set to 350 EUR/kW. 26 Based on Scenario 7, with all other parameters unchanged, the capital cost of RD2 is set to 100 EUR/kW.
Figure 4: Cost Comparison and Intergration of SBSP with Other Technologies in Scenario 1. This figure illustrates a comparative analysis of SBSP with other renewable energy technologies in terms of generation costs, along with the installed capacity of each energy type. Capital costs represent the total annualized investment required for power generation assets over their operational lifetimes, accounting for both initial and operational expenses.
The SBSP-2050 scenarios - Scenario 2-10
In this section, we present findings from the integration of SBSP into the projected 2050 European energy network, evaluating its value and impacts across cost, temporal, and spatial dimensions. The analysis is conducted under Scenario 2-10 to address uncertainties in future energy costs of SBSP.
@@
Fig. 5 provides an overview of the installed capacity, electricity generation, and energy storage utilization of various energy sources under different capital cost scenarios for 2050. The capital costs of RD1 and RD2 are reduced to 6.4–9.7 times and 9.5–14.3 times of solar power in 2050.
@@
This substantial reduction of costs enhances the competitiveness and viability of RD1 in the 2050 scenarios. With RD1, terrestrial wind and solar capacities shrink by up to 75%, while hydropower remains largely unchanged due to its flexibility and reservoir benefits. Due to RD1’s lower generation costs and higher reliability, the generation from solar and wind decreases by 70-87% and 29-81%, respectively. RD1 outperforms both solar and wind across all scenarios, emerging as the dominant electricity generation source. Although battery usage decreases by 59–78%, long-duration storage (notably PHS and hydrogen) remains largely constant, indicating RD1 can handle short-term imbalances yet still depends on seasonal storage solutions.
@@
By contrast, RD2 never appears in the cost-optimal mix under NASA’s 2050 assumptions, suggesting it is not yet competitive with terrestrial renewables or RD1’s more continuous power profile. Collectively, these findings underscore how a substantial reduction in SBSP capital costs can reshape future energy systems, with RD1 emerging as a potentially dominant baseload resource while RD2 lags behind in economic viability under the NASA’s current predictions.
Figure 5: Technology-specific electricity generation and storage supply across scenarios in 2050. Panels (a) and (b) show electricity generation by technology under scenarios with RD1 and RD2, highlighting input data for capital costs and optimized values for installed capacity and generation. Panel (c) presents the total annual supply of each storage technology (PHS, Battery, and Hydrogen) as well as the combined total supply across the same scenarios. PHS pumped hydro storage.
Accounting all system changes together in the system cost, under NASA’s projected 2050 parameters, RD1 (near-baseload SBSP) can reduce total system costs by 7–15% compared to a no-SBSP scenario (Fig. 6). The stable power supply displaces significant portions of intermittent renewables such as wind, solar, and biomass, while also reducing the need for battery storage and long-distance power transmission. Hydropower and nuclear are least affected by SBSP due to their geographical constraints. Total system costs remain unchanged when integrating RD2, due to its absence.
Figure 6: Total system costs across scenarios in 2050. This figure shows the total system cost variations under Scenario 2 to 10 (denoted as S2 to S10). Total system costs include annual capital and variable costs (i.e., fuel, operation, and maintenance) for electricity generation, storage, and transmission.

### v2

@@ Abstract
Meeting net-zero targets remains formidable as terrestrial renewables grapple with intermittency and regional variability. Here, we integrate space-based solar power—a potential near-constant, orbital solar technology—into a high-resolution, Europe-wide capacity-expansion and dispatch model to quantify its contribution under net-zero constraints. We examine two advanced space solar designs: (1) a near-baseload, low technology readiness level concept (heliostat-based representative design one) and (2) a partially intermittent, higher technology readiness level concept (planar-based representative design two), both drawing on NASA’s 2050 cost and performance projections. Our results show that heliostat-based design can reduce total system costs by 7–15%, displace up to 80% of intermittent wind and solar, and cut battery usage by over 70%, if it meets its forecast cost reductions—though long-duration storage (e.g., hydrogen) remains essential for seasonal balancing. By contrast, planar-based design is economically unattractive at its projected 2050 costs. Through extensive sensitivity analyses, we identify cost thresholds at which space-based solar power shifts from cost-prohibitive to complementary and ultimately to a dominant baseload technology. Specifically, heliostat-based design becomes complementary at roughly 14× and dominant at 9× the 2050 solar PV capital cost, benefiting from its continuous power generation. Meanwhile, planar-based design must achieve even lower cost levels (9× to be complementary and 6× to dominate) and would rely on short-duration storage to mitigate its partial intermittency. These findings provide quantified techno-economic benchmarks and reveal alternative net-zero pathways, offering critical guidance for policymakers and industry stakeholders seeking large-scale, centrally coordinated renewable solutions with non- or low-intermittency.
@@ 1 Background
Space-based solar power (SBSP) could offer a centralized, weather-independent energy resource to help address these challenges [6]. By operating above the atmosphere and outside the day-night cycle, SBSP promises continuous gigawatt-scale power. Recent technological milestones suggest it may evolve from a niche concept to a techncially viable solution by the 2030s: multi-junction and lightweight PV cells achieving near 47% efficiency [7, 8], modular in-orbit assembly [9], and successful wireless power demonstrations [10, 11] have all reached mid-range Technology Readiness Levels (TRL 4–5). Meanwhile, reusable launch vehicles have significantly reduced launch costs, with some estimates projecting a Levelized Cost of Energy (LCOE) for SBSP of \$30–80 MWh-1 by 2050 [12, 13, 14, 15]. In parallel, advances in integrated uncertain optimal design for truss configurations [16], interval Riccati equation-based dynamic analysis [17], convex set reliability-based optimal attitude control [18], and multi-objective optimization for robust attitude determination [19] have further strengthened the technical foundations of SBSP. Although considerable uncertainties remain—from in-orbit manufacturing to policy frameworks—major space agencies (NASA, ESA, JAXA) are actively shaping regulatory pathways, motivating the need to understand SBSP’s potential contribution to net-zero goals [14, 20, 21].
@@
  • Integrate SBSP into a continental-scale model: We embed SBSP into a European-scale capacity-expansion and dispatch framework, focusing on meeting electricity demand as defined by ENTSO-E’s 2050 native demand [22], and evaluate its interactions with terrestrial renewables, storage, and transmission infrastructure under net-zero constraints.
  • Assess grid balancing benefits: We quantify how continuous SBSP can reduce storage requirements and transmission investments by offsetting the temporal and spatial variability of terrestrial wind and solar.
@@ 2.1 SBSP Concept
Figure 1(a) illustrates the main stages of an SBSP system, from satellite deployment in orbit to grid integration on Earth. Most designs target geostationary Earth orbit (GEO), where continuous solar exposure enables stable, high-capacity power generation with minimal eclipse times [23]. After launch, robotic systems assemble modular components, such as lightweight photovoltaic (PV) arrays or concentrating mirrors, for solar collection [24, 25]. The harvested electricity is then converted into microwave or laser beams for transmission. Microwave frequencies (1–10 GHz), especially 2.45 GHz or 5.8 GHz [26], are commonly selected to balance transmission efficiency, atmospheric attenuation, and safety constraints [27, 28].
@@
On the ground, rectifying antennas (rectennas) spanning several square kilometers capture the transmitted energy and convert it to direct current (DC), which is then inverted to alternating current (AC) and fed into the existing grid infrastructure. Lightweight rectenna designs allow partial land co-use, and multiple stations can share power from a single satellite [23]. A centralized control center manages satellite pointing, beam transmission, and diagnostics, ensuring real-time coordination and leveraging GEO’s continuous solar availability to offer a high-capacity, dispatchable renewable energy source.
Figure 1: Overview of SBSP Operational Process and System Architectures. (a) Stepwise operational process of a SBSP system, including: (1) launch and installation in space, (2) solar energy collection, (3) conversion to electricity and then to microwave, (4) transmission to Earth, (5) reception and reconversion on the ground, and (6) grid delivery. (b) Main structures of two representative SBSP system designs. The left panel shows the Innovative Heliostat Swarm concept, which is broadly derived from the Alpha Mark III architecture [29]. This design employs reflectors and a central concentrator to continuously focus sunlight throughout the day, using independently operating hexagonal modules arranged in a beehive-like configuration. The right panel depicts the mature Planar Array system with a sandwich architecture: solar collection on one side and microwave transmission on the other. Identical power modules convert solar energy to microwave power and are wirelessly controlled by a central bus, while Earth-facing antennas maintain orientation via gravity gradient forces in GEO.
@@
Although originally conceived by Peter Glaser in 1968 [30], SBSP remained impractical in early evaluations due to prohibitively high costs for launch, in-orbit assembly, and maintenance [31, 32]. Over the past three decades, however, significant advances in solar cell efficiency, wireless power transmission, and reusable launch vehicles have reignited interest in SBSP. A variety of novel designs have since emerged, including the Large Array of Solar Panels [30], Sun Tower Design [32], Modular Symmetrical Concentrator [33], Solar Power Satellite via Arbitrarily Large Phased Array [34], Space-Based Reflectors [35], Inflatable Thin-Film Structures [36], Hypermodular SBSP Design [37], and Tethered Satellite System [38]. Each concept seeks to optimize power density, structural mass, and reliability under the demanding conditions of orbit (Details in the Supplementary Information Section 4).
@@
  • Representative Design One (RD1), the Innovative Heliostat Swarm: Based on the SPS-ALPHA Mark III concept [40, 29, 41], this lower-TRL design uses mirror-like reflectors (heliostats) to direct sunlight to a central concentrator, enabling nearly 99.7% annual power availability [40]. Key innovations include retro-directive RF transmission arrays, high-efficiency PV cells, lightweight modular structures, and autonomous in-orbit deployment. Figure 1(b) shows how reflectors continuously adjust orientation, maximizing solar capture throughout the orbital cycle.
  • Representative Design Two (RD2), the Mature Planar Array: Adapted from JAXA’s Tethered Solar Power Satellite [25] and Caltech’s SSPP design [42], planar design (RD2) includes planar panels whose lower surface faces Earth under gravity-gradient forces. While solar incidence on the upper and lower surfaces varies with orbital geometry, the system achieves roughly 60% annual power availability [25]. Figure 1(b) illustrates its higher TRL, given demonstrated hardware and well-documented performance characteristics.
These two concepts differ in TRL and technical approach yet both incorporate high-efficiency solar capture, wireless power transmission, and modular architectures designed to lower launch mass and total costs. By examining heliostat design (RD1, lower TRL) and planar design (RD2, higher TRL) side-by-side, we provide a balanced assessment of SBSP performance, feasibility, and cost. This approach captures essential trade-offs in orbital assembly complexity, antenna design, and seasonal power availability, ensuring our analysis thoroughly reflects the technical and economic considerations that will guide future SBSP deployment.
@@ 2.2 SBSP Modeling
To assess the system-level impacts of SBSP, we modeled the generation profiles and lifetime costs of NASA’s two representative designs, heliostat design (RD1) and planar design (RD2), for both current (2020) and future (2050) scenarios. Drawing on published technical data [39], we derived hourly generation outputs by accounting for orbital geometry, eclipse periods, solar incidence angles, and design-specific operational constraints (e.g., RD1’s heliostat-based approach vs. RD2’s planar configuration). We assume the geostationary orbit altitude ($\sim 35{,}786\text{ km}$) is negligible relative to the Earth–Sun distance ($\sim 1.49\times 10^{8}\text{ km}$) [43, 39]. Complete derivations and assumptions, including the inverse-square law for irradiance and eclipse scheduling (detailed in the Supplementary Information Section 1).
Generation Profiles. Both heliostat design (RD1) and planar design (RD2) capture solar energy in GEO nearly year-round, but differ in achievable annual availability: heliostat design uses heliostats to concentrate sunlight, enabling up to 99.7% power availability, while planar design’s planar panels achieve approximately 60% due to geometric constraints [40, 25]. Compared to terrestrial solar panels whose power availability is usually 15-30% [44], both heliostat design and planar design show improved power availability. We simulate and compare the hour-by-hour power output per square meter of heliostat design, planar design, and terrestrial solar panels, based on representative orbital parameters, solar irradiation, panel characteristics, and efficiencies (SI, Section 1). Using 2020 as a case study, the results shown in Figure 2 demonstrate the potential advantages of SBSP power generation. The power availability of heliostat design and planar design exceeds that of terrestrial solar by 70% and 42%, respectively. Furthermore, their power output is more continuous, significantly improving efficiency, especially for heliostat design. During the continuous phase, the power output per square meter of heliostat design and planar design is higher at different times compared to terrestrial solar. Over the course of 2020, the power output per square meter of heliostat design and planar design panels is 384% and 233% higher than that of terrestrial solar, respectively. Terrestrial solar has higher output in summer than in winter, while SBSP exhibits the opposite trend, providing complementary power generation in the future. Additionally, the power output of SBSP is significantly less affected by seasonal changes compared to terrestrial solar, indicating its potential for more stable power generation with smaller fluctuations.
@@
Figure 2: Comparison of the variations in power output per square meter of heliostat design (RD1), planar design (RD2), and terrestrial solar panels in 2020. (a) To clarify the power generation curves of heliostat design, planar design, and terrestrial solar, the hourly data were averaged to obtain daily mean values, enabling the observation of overall trends throughout 2020. (b) Hourly data from July 1 to July 15 represent the summer period. (c) Hourly data from December 1 to December 15 represent the winter period.
Cost Analysis. We adopt NASA’s five-phase concept of operations (ConOps) framework (Develop, Assemble, Operate, Maintain, and Dispose) to estimate capital expenditures (CapEx) and fixed operation and maintenance costs (FOM) [39]. The Develop and Assemble phases, covering technology R&D, hardware manufacturing, launch, and in-orbit assembly, are aggregated into CapEx, while Operate, Maintain, and Dispose phases contribute to FOM. Key cost drivers include reusable launch vehicles, modular assembly robotics, and lifetime maintenance. Variable Operations and Maintenance (VOM) costs typically apply to electricity fuel sources. However, since SBSP does not rely on fuel as a direct input for electricity production, this cost category, along with Fuel Costs, is excluded for all renewable electricity production technologies [39]. Under baseline assumptions, we find that near-term (2020) SBSP costs are high, but can drop by 2050 with NASA’s projected efficiencies, launch cost reductions, and technological improvements (SI, Section 2). Building on this foundation, we conducted multiple variable sensitivity analyses, incorporating predicted data for various indicators based on NASA’s projections for 2050 [39]. The estimated costs of heliostat design and planar design in 2020 (the baseline case) and 2050 are shown in Table 1. CapEx refers to the capital expenditures required to acquire, upgrade, or maintain long-term assets, such as equipment, while FOM represents the annual costs associated with the fixed operation and maintenance of these assets. VOM, on the other hand, refers to the variable operation and maintenance costs that change with energy production or system usage [45].
@@
Table 1: Key Parameters [39] and Costs in models (PyPSA-Eur) for SBSP under baseline and 2050 scenarios.
@@
Heliostat Baseline 30 2,028.79 99.7 3
Planar Baseline 30 2,021.95 60 3
Heliostat 2050 30 2,028.79 99.7 3
Planar 2050 30 2,021.95 60 3
[…]
41,071.35 2,806.24 4,901.65 0
62,335.91 4,506.15 7,687.48 0
2,915.31 119.15 267.87 0
3,701.30 207.71 396.59 0
@@ 3 Results
Interaction with terrestrial renewables in 2050 power systems
We individually integrate heliostat design (RD1) and planar design (RD2) into the PyPSA-Eur model for minimum-cost optimization under both current (2020) and future (2050) scenarios (see “Methods”). The results for the current scenario are provided in SI Section 5, as neither heliostat design nor planar design is selected in the optimized system. This outcome highlights the economic infeasibility of SBSP under present-day conditions due to its high capital cost. We evaluate the value and impacts of SBSP across cost, temporal, and spatial dimensions under three main 2050 scenarios. These scenarios—termed low, medium, and high SBSP market opportunity—differ in the capital costs of available generation resources, which include solar photovoltaics (PV), onshore and offshore wind, reservoir hydro, biomass, run-of-river hydro, and nuclear. The capital costs of these resources, listed in Table 2, are lowest in the low-opportunity scenario and highest in the high-opportunity scenario.
[…]
Low Medium High
Solar 27.68 32.29 41.74
Onshore wind 72.36 103.75 121.46
Offshore wind 197.30 211.70 222.12
Reservoir 216.67 216.67 216.67
Biomass 226.31 229.89 233.40
Run-of-river 318.33 319.88 321.43
Nuclear 443.76 443.76 443.76
@@
Fig. 3 provides an overview of the installed capacity and electricity generation of various energy sources under different capital cost scenarios for 2050. The capital costs of heliostat design (RD1) and planar design (RD2) are reduced to 6.4–9.7 times and 9.5–14.3 times of solar power in 2050.
Figure 3: Technology-specific installed capacity and electricity generation across scenarios in 2050. Panels (a) and (b) show installed capacity and electricity generation by technology under 3 market opportunity scenarios with heliostat design (RD1) and planar design (RD2).
@@
This substantial reduction of costs enhances the competitiveness and viability of heliostat design (RD1) in the 2050 scenarios. With heliostat design, terrestrial wind and solar capacities shrink by up to 75%, while hydropower remains largely unchanged due to capacity constraints and the assumption of using existing capacity in Europe. Due to heliostat design’s lower generation costs and higher reliability, the generation from solar and wind decreases by 70-87% and 29-81%, respectively. Heliostat design outperforms both solar and wind across all scenarios, emerging as the dominant electricity generation source.
@@
By contrast, planar design (RD2) never appears in the cost-optimal mix under NASA’s 2050 assumptions, suggesting it is not yet competitive with terrestrial renewables or heliostat design’s more continuous power profile. Collectively, these findings underscore how a substantial reduction in SBSP capital costs can reshape future energy systems, with heliostat design emerging as a potentially dominant baseload resource while planar design lags behind in economic viability under the NASA’s current predictions.